OpenStudy (anonymous):

I'm in the midst of solving the heat equation with boundary conditions. My question is that I'm left with what looks like a Fourier Sine Series but I'm not sure. It's in the format sin[(2n+1)*pi*x/(2L)] as opposed to the normal sin(n*pi*x/L) Can I still use the same method of finding the coefficient?

OpenStudy (anonymous):

u have done some mistake so far

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

there is nothing wrong with that u must have boundary condition like this to get that \[0\le x \le L\\ T(x=0)=0 \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{\partial T }{\partial x } (x=L)=0\] method is the same

OpenStudy (anonymous):

Those are the boundary conditions. I just wasn't sure if setting up the integral was the same. Thanks @mukushla